OpenStudy (anonymous):

the length of a rectangle is 5 m more than twice its width ,and the area of the rectangle is 88m^2 . find the dimensions . length : width :

OpenStudy (lgbasallote):

l = 5 + 2w w = w a = lw a = (5+2w)(w)

OpenStudy (anonymous):

a=2w^2+5w

OpenStudy (lgbasallote):

now equate that to your area then use completing square/quadratic formula

OpenStudy (anonymous):

i dont know how to use it

OpenStudy (lgbasallote):

2w^2 + 5w = 88 2w^2 + 5w - 88 = 0 a = 2; b = 5; c = -8 quadratic form: x = -b +/- sqrt (b^2 - 4ac) all over 2a

OpenStudy (anonymous):

i just need the answer of the length and the width !!

OpenStudy (lgbasallote):

then you better start substituting values ^_^

Directrix (directrix):

Continuing from above: 2w^2 + 5w - 88 = 0 will factor (w+8)(2w-11) = 0 Zero Product Property w +8 = 0 or 2w -11 = 0 w = -8 or 2w = 11 w = -8 (discard for this application) or w = 11/2 = 5.5 w = 5.5 meters and Length = 2(5.5) + 5 = 16 meters